You have two moles of gaseous carbon dioxide at room temperature to which you add three kilocalories of heat. a. What is its increase in temperature? b. What is its change in internal energy?
Answer #1


internal energy E1 = (3/2)*n*R*T1


energy added = 3000 calorie = 3000*4.2 = 12600 J


total energy E2 = (3/2)*n*R*T2

E2 = E1 + 12600


E2 - E1 = 12600

(3/2)*2*8.314*dT = 12600


dT = 505 K


==================

change = 12600 J